An object of mass 4kg moves in a circle at radius ,8m at uniform speed of 32m/s
(1) calculate the angular velocity
1
Expert's answer
2018-11-26T16:13:10-0500

m=4kg

R=8m

v=32 m/s

Angular velocity = linear velocity/Radius

ω=v/R=(32 m⁄s)/(8 m)=4 rad⁄s

if the body weight is given, then you probably need to find the centrifugal force

F=m v^2/R=mωv=4∙4∙32=512(N)